Find (s,∘,w)(1)for s(x)=4x2+3x−3and w(x)=3x2+x−4.
(s,∘,w)(1)=

Answer by greenestamps(13198)



w(1) = 3(1^2)+1-4 = 0

s(w(1)) = s(0) = 0+0-3 = -3